OpenStudy (anonymous):

how do i change x^2-8x = M to a vertex form to find where the minimum occurs at, i guess the x and y. :D

OpenStudy (anonymous):

Well, you need to complete the square to get it into a form where you can easily spot the vertex. y=x^2-8x y=x^2-8x+16-16 y=(x-4)^2-16 So your vertex is (4,-16)